# UV-Attack: Physical-World Adversarial Attacks for Person Detection via Dynamic-NeRF-based UV Mapping

### v1

@@ 4.3 Physical Attack Results
Figure 3: Physical attack success rates with different detectors and confidence thresholds. The attack is trained on FastRCNN model, and the IoU threshold is set as 0.5.
@@ 4.4 Ablation Study
Render time for one image $\sim$10h $\sim$2h $\sim$15min <0.5s
@@ A.1 Attack results on more object detection models
To evaluate the transferability of our attack, we trained the adversarial textures on FastRCNN (Girshick, 2015) and tested their transferability across a variety of person detectors with diverse backbones. These include models such as Deformable DETR (Zhu et al., 2020), RetinaNet (Lin et al., 2018), SSD (Liu et al., 2016), FCOS (Tian et al., 2019), as well as more recent detectors like RTMDET (Lyu et al., 2022), YOLOv8 (Jocher et al., 2023), ViTDET (Li et al., 2022), and Co-DETR (Zong et al., 2023). Notably, Co-DETR achieves the highest mean Average Precision (mAP) of 66% on the COCO object detection task. ViTDET, which employs the Vision Transformer as its backbone, allows us to further assess the attack’s effectiveness across different backbone architectures. Co-DETR also supports different backbones. Our experiment tests Co-DETR on two different backbones, ResNet50 and Swin Transformer. We implement the Co-DETR and ViTDET using MMDET toolbox (Chen et al., 2019).
@@
$\text{ASR}=\frac{1}{|\mathcal{D}|}\sum_{i\in\mathcal{D}}\mathbb{I}\left(\exists b_{i}\in B_{i}:\text{IoU}(b_{i},g_{i})>\tau_{\text{IoU}}\land\text{Conf}(b_{i})>\tau_{\text{Conf}}\land\text{label}(b_{i})=\text{person}\right)$ (6)
@@ B.1 The influence of environments
To evaluate the attack performance in different environments, we test the attack success rate and mAP on a new indoor dataset (Quattoni & Torralba, 2009), which contains 67 Indoor categories, and a total of 15620 images. The number of images varies across categories, but there are at least 100 images per category. We select 13 common scenarios: bedroom, airport, office, etc. The attack results are shown in Figure 10. It is shown that all scenarios achieve a high ASR when the confidence threshold is set as 0.5. For the mAP, all scenarios are less than 0.2 and 6 scenarios are less than 0.1, including bedroom, bathroom, airport, office, library, and church.
@@ Appendix C Additional experiments in the physical environments
We demonstrate the attack effects in three different scenarios: a hallway, a lawn, and a library. For each scenario, we recorded a video of approximately 1 minute with a frame rate of 30 FPS, using an iPhone 13 as the recording device. The confidence threshold is set as 0.5. The results are shown in Figure 16. The adversarial patch is trained on the Faster-RCNN model. In these three scenarios, we achieve ASR of 0.84, 0.83, and 0.80 respectively. The experiments show that we can successfully fool Faster-RCNN in various scenarios.
@@
Figure 20: Comparisons of transferability with the naturalness constraints. Here we add the classifier-free guidance to generate natural images. It is shown that our method outperforms previous SOTA methods regarding transferability on Retina, SSD, and DETR and shows competitive results on YOLOv3 and FCOS.

### v2

@@ 4.3 Physical Attack Results
Figure 3: Physical attack success rates with different detectors and confidence thresholds. The attack is trained on the FastRCNN model, and the IoU threshold is set as 0.5.
@@ 4.4 Ablation Study
Render time for one image $\sim$5h $\sim$2h $\sim$15min <0.5s
@@ 5 Conclusion
Acknowledge
This work was supported in part by HK RGC GRF under Grant PolyU 15201323. Additionally, we would like to thank Wenxuan Zhang at Hong Kong Polytechnic University for his assistance with the ablation experiments.
@@ A.1 Attack results on more object detection models
To evaluate the transferability of our attack, we trained the adversarial textures on FastRCNN (Girshick, 2015) and tested their transferability across a variety of person detectors with diverse backbones. These include models such as Deformable DETR (Zhu et al., 2020), RetinaNet (Lin et al., 2018), SSD (Liu et al., 2016), FCOS (Tian et al., 2019), as well as more recent detectors like RTMDET (Lyu et al., 2022), YOLOv8 (Jocher et al., 2023), ViTDET (Li et al., 2022), and Co-DETR (Zong et al., 2023). Notably, Co-DETR achieves the highest mean Average Precision (mAP) of 66% on the COCO object detection task. ViTDET, which employs the Vision Transformer as its backbone, allows us to further assess the attack’s effectiveness across different backbone architectures. Co-DETR also supports different backbones. Our experiment tests Co-DETR on two different backbones, ResNet50 and Swin Transformer. We implement Co-DETR and ViTDET using the MMDET toolbox (Chen et al., 2019).
@@
$\text{ASR}=1-\frac{1}{|\mathcal{D}|}\sum_{i\in\mathcal{D}}\mathbb{I}\left(\exists b_{i}\in B_{i}:\text{IoU}(b_{i},g_{i})>\tau_{\text{IoU}}\land\text{Conf}(b_{i})>\tau_{\text{Conf}}\land\text{label}(b_{i})=\text{person}\right)$ (6)
@@ B.1 The influence of environments
To evaluate the attack performance in different environments, we test the attack success rate and mAP on a new indoor dataset (Quattoni & Torralba, 2009), which contains 67 Indoor categories, and a total of 15620 images. The number of images varies across categories, but there are at least 100 images per category. We select 13 common scenarios: bedroom, airport, office, etc. The attack results are shown in Figure 10. It is shown that all scenarios achieve a high ASR when the confidence threshold is set as 0.5. For the mAP, all scenarios are less than 0.2, and 6 scenarios are less than 0.1, including bedroom, bathroom, airport, office, library, and church.
@@ Appendix C Additional experiments in the physical environments
We demonstrate the attack effects in three different scenarios: a hallway, a lawn, and a library. For each scenario, we recorded a video of approximately 1 minute with a frame rate of 30 FPS, using an iPhone 13 as the recording device. The confidence threshold is set as 0.5. The results are shown in Figure 16. The adversarial patch is trained on the Faster-RCNN model. In these three scenarios, we achieve ASR of 0.84, 0.83, and 0.80, respectively. The experiments show that we can successfully fool Faster-RCNN in various scenarios.
@@
Figure 20: Comparisons of transferability with the naturalness constraints. Here, we add the classifier-free guidance to generate natural images. It is shown that our method outperforms previous SOTA methods regarding transferability on Retina, SSD, and DETR and shows competitive results on YOLOv3 and FCOS.